OpenStudy (anonymous):

Question 1.1. Which property is demonstrated? 7 + 9 + 3 = 7 + 3 + 9 associative property of addition associative property of multiplication commutative property of addition commutative property of multiplication Question 2.2. Simplify 10 • 4 • 4 • 50 800 2000 4000 8000 Question 3.3. Which property is demonstrated? 72 + (28 + 93) = (72 + 28) + 93 associative property of addition associative property of multiplication commutative property of addition commutative

OpenStudy (anonymous):

@satellite73 help??

OpenStudy (shinalcantara):

since it doesn't involve groupings using parenthesis and it's just the arrangement then it is commutative property of addition

OpenStudy (anonymous):

\[7 + 9 + 3 = 7 + 3 + 9\] numbers change places, commutative property

OpenStudy (anonymous):

whao \[10\times 4\times 4\times 50=8000\] http://www.wolframalpha.com/input/?i=10*4*4*50

OpenStudy (anonymous):

\[72 + (28 + 93) = (72 + 28) + 93 \] parentheses move, associative property